on weekends, as part of his exercise routine, Carl goes for a run, partly on paved trails and partly across rough terrain. He runs at 10km/h on the trails, but his speed is reduced to 5km/h on the rough terrain. One day, he ran 12km in 1.5h. How far did he run on the rough terrain?
Answer by richwmiller(17219)

5*a+10*b=12,
a+b=1.5
a=0.6 hr=36 minutes b=0.9 hr=54 minutes
5*.6=3 km on rough
.9*10=9 km on paved